OpenStudy (anonymous):

WHATS THE Y INTER OF 19Y+20X=5 ?

OpenStudy (anonymous):

let x = 0 19y + 20(x) = 5 now solve for y 19y = 5 or y = 19/5

OpenStudy (mathmate):

y=5/19